Find the coordinates of the midpoint of line HK if H (-1, 2) and K (-7, -4).
 
 

Grade:9

1 Answers

Tamanna Grover
3 years ago
Hi friend, in these kind of questions, you are to apply the formula: required x coordinate=(X1+X2)/2 and required y coordinate=(y1+y2)/2. Therefore, we get {(-1-7)/2},{(2-4)/2} which is (-4,-1).